STRUCTURE DETERMINATION OF TWO STEREOISOMERS OF SEVOFLURANE DIMER BY CHIRPED PULSE FTMW SPECTROSCOPY

\underline{NATHAN A. SEIFERT}, CRISTOBAL PEREZ, DANIEL P. ZALESKI, JUSTIN L. NEILL, BROOKS H. PATE, Department of Chemistry, University of Virginia, McCormick Rd., Charlottesville, VA 22904-4319; ALBERTO LESARRI, MONTSERRAT VALLEJO, Departamento de Química Física y Química Inorgánica, Facultad de Ciencias, Universidad de Valladolid, E-47011 Valladolid, Spain; EMILIO J. COCINERO, FERNANDO CASTANO, Departamento de Química Física, Facultad de Ciencia y Tecnología, Universidad del País Vasco (UPV/EHU), Campus de Leioa, Ap. 644, E-48080 Bilbao, Spain.

Two stereoisomers of sevoflurane dimer have been detected using chirped pulse FTMW spectroscopy from 2-8 GHz. The identified complexes are distinguished by their differing helicities, and together both isomers form a diastereomeric pair, one being homochiral (RR/SS) and the other heterochiral (RS/SR). For both isomers, all 8 ^{13}C isotopologues have been assigned, and two ^{18}O isotopologues have also been detected for the homochiral isomer, for a total of 18 isotopologues. MP2/6-311++g(d,p) calculations predict the heterochiral isomer as 1.2 kJ mol^{−1} above the homochiral species, which is consistent with the observed relative intensities between the two species. A summary of these microwave results, including a comparison between the Kraitchman and ab initio structures, will be presented.